Trevor, currently 22 years old, is half of his dad's age six years ago. How old is Trevor's dad?
Answer by ankor@dixie-net.com(22740)

revor, currently 22 years old, is half of his dad's age six years ago.
How old is Trevor's dad?
:
Let d = Dad's present age
:
Just write an equation for what it says:
.5(d-6) = 22
.5d - 3 = 22
.5d = 22 + 3
.5d = 25
multiply both sides by 2
d = 50 yrs old is Dad's present age